# Error Propagation Mechanisms and Compensation Strategies for Quantized Diffusion Models

### v1

@@ Abstract
Diffusion models have transformed image synthesis by establishing unprecedented quality and creativity benchmarks. Nevertheless, their large-scale deployment faces challenges due to computationally intensive iterative denoising processes. Although post-training quantization(PTQ) provides an effective pathway for accelerating sampling, the iterative nature of diffusion models causes stepwise quantization errors to accumulate progressively during generation, inevitably compromising output fidelity. To address this challenge, we develop a theoretical framework that mathematically formulates error propagation in Diffusion Models (DMs), deriving per-step quantization error propagation equations and establishing the first closed-form solution for cumulative error. Building on this theoretical foundation, we propose a timestep-aware cumulative error compensation scheme. Extensive experiments across multiple image datasets demonstrate that our compensation strategy effectively mitigates error propagation, significantly enhancing existing PTQ methods to achieve state-of-the-art(SOTA) performance on low-precision diffusion models.
@@ 1 Introduction
DMspodell2023sdxl chen2024pixart have established themselves as highly effective deep generative frameworks across diverse domains, including image synthesisho2020denoising , video creationyang2024cogvideox and image translationsasaki2021unit etc. Compared with conventional SOTA generative adversarial networks (GANs), DMs exhibit superior stability, free from the common pitfalls of model collapse and posterior collapse, which ensures more reliable and diverse output generation. Although diffusion models demonstrate remarkable capabilities in generating high-fidelity and diverse images, their substantial computational and memory overhead hinders widespread adoption. This complexity primarily stems from two factors: first, the reliance on complex deep neural networks (DNNs) for noise estimation; second, the requirement for an iterative progressive denoising process to maintain synthesis quality, which may involve up to 1,000 iterative steps, substantially increasing computational demands.
@@
To address the substantial computational demands of noise estimation in diffusion models, researchers employ quantization techniques to accelerate inference across all denoising steps. Depending on whether they require fine-tuning, quantization methods can be categorized into Quantization-Aware Training (QAT)wu2020easyquant and PTQcai2020zeroq . QAT necessitates retraining neural networks with simulated quantization and hyperparameter optimization, which introduces significant computational overhead and deployment complexity, making it unsuitable for compute-intensive diffusion model training. In contrast, PTQ directly derives quantization correction coefficients through calibration data statistics, thus avoiding the high-cost retraining process inherent to diffusion models. Although PTQ has been widely studied in traditional vision tasks such as image classification and object detectionbhalgat2020lsq+ , it faces many different challenges in diffusion models. The architectural characteristics and training strategies of diffusion models inherently lead to the widespread presence of outliers in weight distributions, while activation values exhibit step-wise distributional variations across time stepsqdiffusion . These properties pose significant challenges for quantization by inducing substantial step-wise error. Furthermore, the iterative denoising mechanism amplifies error accumulation across successive steps, where quantization errors progressively accumulate during the sampling trajectory, ultimately degrading generation fidelity. Recent advancements in PTQwu2024ptq4dit qdiffusion mixdq have predominantly focused on minimizing quantization errors at individual denoising steps. However, these approaches systematically neglect the critical analysis of error propagation dynamics throughout the iterative sampling trajectory. Consequently, current solutionssvdqunat remain constrained to 4-bit quantization of both weights and activations while maintaining acceptable quality degradation, as cumulative errors across sequential denoising stages fundamentally limit lower-bit quantization viability.
@@
Contrarily, our work focuses on the quantization error propagation problem in diffusion models and proposes the first timestep-aware cumulative error compensation scheme, called TCEC. First, we construct an error propagation equation by taking the DDIMsong2020denoising as a paradigmatic case, presenting the field’s inaugural closed-form solution for cumulative error. However, direct computation of cumulative errors proves computationally infeasible. Subsequently, through the implementation of timestep-sensitive online rapid estimation for per-step quantization error, we achieve a notable simplification of the computational complexity inherent in cumulative error modeling. Finally, as shown in Fig.1, we incorporate a cumulative error correction term in each generation step, dynamically mitigating errors induced by quantization. In summary, our contributions are:
@@
Notably, TCEC maintains orthogonality to existing state-of-the-art PTQ algorithmssvdqunat vidit that minimize per-step quantization errors. Additionally, TCEC originates from rigorous theoretical derivation, and while its derivation process uses DDIM as an example, it is equally applicable to other solvers such as DPM++lu2022dpm .
@@ 2.1 Diffusion Model
Diffusion models are a family of probabilistic generative models that progressively destruct real data by injecting noise, then learn to reverse this process for generation, represented notably by denosing diffusion probabilistic models(DDPMs)ho2020denoising . DDPM is composed of two Markov chains of $T$ steps. One is the forward process, which incrementally adds Gaussian noises into real sample $x_{0}\in q(x_{0})$, In this process, a sequence of latent variables $x_{1:T}=[x_{1},x_{2},...,x_{T}]$ are generated in order and the last one $x_{T}$ will approximately follow a standard Gaussian:
@@
where $p(\mathbf{x}_{T})\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, with $\mu_{\theta}(\mathbf{x}_{t},t)$ denotes the noise estimation model, and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ denotes the variance for sampling which can be fixed to constantsluo2022understanding . The denoising process, constrained by the Markov chain, requires a huge number of iterative time steps in DDPM. DDIM generalizes the diffusion process to non-Markovian processes, simulating the diffusion process with fewer steps. It has replaced DDPM as the mainstream inference strategy. Our work focuses on accelerating the inference of the noise estimation model in DDIM, with a training-free PTQ process.
@@ 2.2 Post-Training Quantization
The noise estimation models such as UNETpodell2023sdxl or Transformerchen2024pixart yang2024cogvideox zheng2024open exhibit high computational complexity, rendering the sampling of diffusion models computationally expensive. PTQ transforms the weights and activations of the full-precision model into a low-bit format, enabling the model’s inference process to utilize the integer matrix multiplication units on the target hardware platform and accelerating the computational processjacob2018quantization zeng2025abq zeng2024gqsa . Prior studies, such as PTQD he2023ptqd and Q-DM li2023q , have explored the application of quantization techniques for diffusion models. Q-Diffusion qdiffusion and PTQ4DM shang2023post first achieved 8-bit quantization in text-to-image generation tasks. Subsequent research refined these methodologies through strategies such as sensitivity analysis yang2023efficient and timestep-aware quantization huang2024tfmq wang2023towards . Among these works, MixDQmixdq introduces metric-decoupled sensitivity analysis and develops an integer programming-based method to derive optimal mixed-precision configurations. Qua2SeDiMoqua2sedimo enables high-quality mixed-precision quantization decisions for a wide range of diffusion models, from foundational U-Nets to state-of-the-art Transformers, extending the quantization lower bounds for image generation tasks to W4A8. SVDQuantsvdqunat enhances quantization performance by integrating fine-grained quantization with singular value decomposition (SVD)-based weight decomposition, achieving W4A4 quantization while maintaining acceptable quality degradation. Meanwhile, ViDiT-Q vidit further explores quantization for video generation tasksyang2024cogvideox zheng2024open , achieving W8A8 and W4A8 with negligible degradation in visual quality and metrics. These works minimize per-step errors through more precise quantization approximations at the layer level, yet overlook the error propagation in the diffusion process. Particularly in video generation tasks, which require a greater number of inference steps, the issue of error propagation becomes exacerbated, and accumulated errors across sequential denoising stages fundamentally constrain the feasibility of lower-bit quantization. Although a few studieschu2024qncd timestep have acknowledged the error propagation issue and attempted solutions, they lack rigorous theoretical analysis. Their experimental designs rely on ad-hoc fitting to specific tasks and are only validated on simple datasets such as CIFAR10.
@@ 3.1 Error Propagation Mechanisms
The iterative denoising process of diffusion models corresponds to the discrete approximation of the probability flow ordinary differential equation(PF-ODE), the noise at each time step $t\in[T,...,1]$ is computed from $x_{t}$ by a full-precision noise estimation model $\mu_{\theta}$ whose weights are fixed at all steps. Based on DDIM-solver, we can calculate the sample $\mathbf{x}_{t-1}$ at time t1t-1 as follows
@@
where $\alpha_{t}$ is a constant related to the noise schedule $\beta_{t}$ and the specific relationship is $\frac{\alpha_{t}}{\alpha_{t}-1}=1-\beta_{t}$. Since $\beta_{t}\in(0,1)$, $\alpha_{t}$ is monotonically decreasing with respect to $t$. We denote the quantized version of the noise estimation model as $\tilde{\mu}_{\theta}$. When the input remains unchanged, it is formulated by
@@
where $\varepsilon_{t}$ means per-step quantization error which is introduced due to model quantization and only relies on the module at iteration $t$ and is independent of the others. The amount of error accumulated by continuously running the first Tt+1T-t+1 denoising steps is called the cumulative error $\delta_{t}$, then the input including cumulative error can be expressed as $\tilde{x}_{t}=x_{t}+\delta_{t}$. Consequently, referring to Eq. 3, the iterative process of the quantized model can be expressed as
@@
Based on Eq. 4 and the definition of cumulative error, Eq.5 can be reformulated as
@@
Applying the first-order Taylor expansion, $\mu_{\theta}(x_{t}+\delta_{t},t)$ is approximated as $\mu_{\theta}(x_{t},t)+\mathbf{J}_{x_{t}}\delta_{t}$. Substituting it into Eq. 6, we can obtain the error propagation equation that relates the per-step quantization error to the cumulative error
@@
in which $A_{t}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}I+B_{t}*\mathbf{J}_{x_{t}}$, $B_{t}=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}$ and $J_{x_{t}}=\nabla_{x_{t}}\mu_{\theta}(x_{t},t)$ is the Jacobian matrix of the denoising model $\mu_{\theta}$. Given that $\delta_{T}=0$, when we recursively expand Eq.7 from $T$ to $t$, the cumulative error $\delta_{t}$ can be derived as
@@ 3.2 Simplify Computational Complexity
Consequently, the inverse of the propagation matrix $\mathbf{A}_{j}$ can be approximately represented as $\frac{\sqrt{\alpha_{j}}}{\sqrt{\alpha_{j-1}}}\mathbf{I}$. By expanding the product terms in Eq. 8 and reducing the intermediate terms, the final result can be expressed as
@@
Plugging Eq. 9 into Eq. 8, we have
@@
Therefore, we can effectively reformulate Eq.10 as
@@
How to determine the value of the parameter $m$? Based on our discussion in Sec.3.1, the cumulative error at step t1t-1 is related to the cumulative error $\delta_{t}$ from steps $T,\ldots,t+1]$ and the per-step quantization error $\varepsilon_{t}$ at step $t$. By substituting the correction term $\Delta_{t}$ into Eq.7, we have
@@
where $\widehat{\delta_{t-1}}$ is the corrected cumulative error and it should exhibits a strictly smaller upper bound in norm compared to $\delta_{t-1}$, signifying a more refined and accurate error representation. Based on this, we can solve for the reasonable value of $m$. Under mild regularity conditions, there exists $\sigma>0$ independent of the timestep $k$, such that the per-step quantization error satisfies $\|\varepsilon_{k}\|\leq\sigma,\forall k$. Based on this condition, we can deduce from $\|\widehat{\delta_{0}}\|\leq\sigma\sum_{t=1}^{T}\left(\prod_{k=1}^{t-1}\rho_{k}\right)\|\mathbf{B}_{t}\|\leq\|\delta_{0}\|$ that $m=1$. We provide a complete proof to this theorem in Appendix 7.1. This implies that the cumulative error at any timestep is only related to the per-step quantization errors of the two immediately preceding steps. Thus, the final cumulative error correction term can be reformulated as
@@
Table 1: Quantization Performance Comparison of Different Models. SDXL and SDXL-Turbo generate at $512^{2}$ resolution, while PixArt achieves $1024^{2}$. Evaluation metrics include FID (distribution distance)fid , IR (human preference)ir , LPIPS (perceptual similarity)lpips , and PSNR (numerical fidelity against 16-bit references)svdqunat .
[…]

### v2

@@ Abstract
Diffusion models have transformed image synthesis by establishing unprecedented quality and creativity benchmarks. Nevertheless, their large-scale deployment faces challenges due to computationally intensive iterative denoising processes. Although post-training quantization (PTQ) provides an effective pathway for accelerating sampling, the iterative nature of diffusion models causes stepwise quantization errors to accumulate progressively during generation, inevitably compromising output fidelity. To address this challenge, we develop a theoretical framework that mathematically formulates error propagation in Diffusion Models (DMs), deriving per-step quantization error propagation equations and establishing the first closed-form solution for cumulative error. Building on this theoretical foundation, we propose a timestep-aware cumulative error compensation scheme. Extensive experiments on multiple image datasets demonstrate that our compensation strategy effectively mitigates error propagation, significantly enhancing existing PTQ methods. Specifically, it achieves a 1.2 PSNR improvement over SVDQuant on SDXL W4A4, while incurring only an additional << 0.5% time overhead.
@@ 1 Introduction
DMs (Podell et al., 2023; Chen et al., 2024) have established themselves as highly effective deep generative frameworks across diverse domains, including image synthesis (Ho et al., 2020), video creation (Yang et al., 2024) and image translation (Sasaki et al., 2021) etc. Compared with conventional SOTA generative adversarial networks (GANs), DMs exhibit superior stability, free from the common pitfalls of model collapse and posterior collapse, which ensures more reliable and diverse output generation. Although diffusion models demonstrate remarkable capabilities in generating high-fidelity and diverse images, their substantial computational and memory overhead hinders widespread adoption. This complexity primarily stems from two factors: first, the reliance on complex deep neural networks (DNNs) for noise estimation; second, the requirement for an iterative progressive denoising process to maintain synthesis quality, which may involve up to 1,000 iterative steps, substantially increasing computational demands.
@@
To address the substantial computational demands of noise estimation in diffusion models, researchers employ quantization techniques to accelerate inference across all denoising steps. Depending on whether they require fine-tuning, quantization methods can be categorized into Quantization-Aware Training (QAT) (Wu et al., 2020) and PTQ (Cai et al., 2020). QAT necessitates retraining neural networks with simulated quantization and hyperparameter optimization, which introduces significant computational overhead and deployment complexity, making it unsuitable for compute-intensive diffusion model training. In contrast, PTQ directly derives quantization correction coefficients through calibration data statistics, thus avoiding the high-cost retraining process inherent to diffusion models. Although PTQ has been widely studied in traditional vision tasks such as image classification and object detection (Bhalgat et al., 2020), it faces many different challenges in diffusion models. The architectural characteristics and training strategies of diffusion models inherently lead to the widespread presence of outliers in weight distributions, while activation values exhibit step-wise distributional variations across time steps (Li et al., 2023a). These properties pose significant challenges for quantization by inducing substantial step-wise error. Furthermore, the iterative denoising mechanism amplifies error accumulation across successive steps, where quantization errors progressively accumulate during the sampling trajectory, ultimately degrading generation fidelity. Recent advancements in PTQ (Wu et al., 2024; Li et al., 2023a; Zhao et al., 2024b) have predominantly focused on minimizing quantization errors at individual denoising steps. However, these approaches systematically neglect the critical analysis of error propagation dynamics throughout the iterative sampling trajectory. Consequently, current solutions (Li et al., 2024b) remain constrained to 4-bit quantization of both weights and activations while maintaining acceptable quality degradation, as cumulative errors across sequential denoising stages fundamentally limit lower-bit quantization viability.
@@
Contrarily, our work focuses on the quantization error propagation problem in diffusion models and proposes the first timestep-aware cumulative error compensation scheme, called TCEC. First, we construct an error propagation equation by taking the DDIM (Song et al., 2020) as a paradigmatic case, presenting the field’s inaugural closed-form solution for cumulative error. However, direct computation of cumulative errors proves computationally infeasible. Subsequently, through the implementation of timestep-sensitive online rapid estimation for per-step quantization error, we achieve a notable simplification of the computational complexity inherent in cumulative error modeling. Finally, as shown in Figure 1, we incorporate a cumulative error correction term in each generation step, dynamically mitigating errors induced by quantization. In summary, our contributions are:
@@
Notably, TCEC maintains orthogonality to existing state-of-the-art PTQ algorithms (Li et al., 2024b; Zhao et al., 2024a) that minimize per-step quantization errors. Additionally, TCEC originates from rigorous theoretical derivation, and while its derivation process uses DDIM as an example, it is equally applicable to other solvers such as DPM++ (Lu et al., 2022).
@@ 2.1 Diffusion Model
Diffusion models are a family of probabilistic generative models that progressively destruct real data by injecting noise, then learn to reverse this process for generation, represented notably by denosing diffusion probabilistic models (DDPMs) (Ho et al., 2020). DDPM is composed of two Markov chains of $T$ steps. One is the forward process, which incrementally adds Gaussian noises into real sample $x_{0}\in q(x_{0})$, In this process, a sequence of latent variables $x_{1:T}=[x_{1},x_{2},...,x_{T}]$ are generated in order and the last one $x_{T}$ will approximately follow a standard Gaussian:
@@
where $p(\mathbf{x}_{T})\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, with $\mu_{\theta}(\mathbf{x}_{t},t)$ denotes the noise estimation model, and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ denotes the variance for sampling which can be fixed to constants (Luo, 2022). The denoising process, constrained by the Markov chain, requires a huge number of iterative time steps in DDPM. DDIM generalizes the diffusion process to non-Markovian processes, simulating the diffusion process with fewer steps. It has replaced DDPM as the mainstream inference strategy. Our work focuses on accelerating the inference of the noise estimation model in DDIM, with a training-free PTQ process.
@@ 2.2 Post-Training Quantization
The noise estimation models such as UNET (Podell et al., 2023) or Transformer (Chen et al., 2024; Yang et al., 2024; Zheng et al., 2024) exhibit high computational complexity, rendering the sampling of diffusion models computationally expensive. PTQ transforms the weights and activations of the full-precision model into a low-bit format, enabling the model’s inference process to utilize the integer matrix multiplication units on the target hardware platform and accelerating the computational process (Jacob et al., 2018). Prior studies, such as PTQD (He et al., 2023) and Q-DM (Li et al., 2023b), have explored the application of quantization techniques for diffusion models. Q-Diffusion (Li et al., 2023a) and PTQ4DM (Shang et al., 2023) first achieved 8-bit quantization in text-to-image generation tasks. Subsequent research refined these methodologies through strategies such as sensitivity analysis (Yang et al., 2023) and timestep-aware quantization (Huang et al., 2024a; Wang et al., 2023). Among these works, MixDQ (Zhao et al., 2024b) introduces metric-decoupled sensitivity analysis and develops an integer programming-based method to derive optimal mixed-precision configurations. Qua2SeDiMo (Mills et al., 2025) enables high-quality mixed-precision quantization decisions for a wide range of diffusion models, from foundational U-Nets to state-of-the-art Transformers, extending the quantization lower bounds for image generation tasks to W4A8. SVDQuant (Li et al., 2024b) enhances quantization performance by integrating fine-grained quantization with singular value decomposition (SVD)-based weight decomposition, achieving W4A4 quantization while maintaining acceptable quality degradation. Meanwhile, ViDiT-Q (Zhao et al., 2024a) further explores quantization for video generation tasks (Yang et al., 2024; Zheng et al., 2024), achieving W8A8 and W4A8 with negligible degradation in visual quality and metrics. These works minimize per-step errors through more precise quantization approximations at the layer level, yet overlook the error propagation in the diffusion process. Particularly in video generation tasks, which require a greater number of inference steps, the issue of error propagation becomes exacerbated, and accumulated errors across sequential denoising stages fundamentally constrain the feasibility of lower-bit quantization. Although a few studies (Chu et al., 2024; Yao et al., 2024) have acknowledged the error propagation issue and attempted solutions, they lack rigorous theoretical analysis. Their experimental designs rely on ad-hoc fitting to specific tasks and are only validated on simple datasets such as CIFAR10.
@@ 3.1 Error Propagation Mechanisms
The iterative denoising process of diffusion models corresponds to the discrete approximation of the probability flow ordinary differential equation(PF-ODE), the noise at each time step $t\in[T,...,1]$ is computed from $x_{t}$ by a full-precision noise estimation model $\mu_{\theta}$ whose weights are fixed at all steps. Based on DDIM-solver, we can calculate the sample $\mathbf{x}_{t-1}$ at time t1t-1 as follows:
@@
where $\alpha_{t}$ is a constant related to the noise schedule $\beta_{t}$ and the specific relationship is $\frac{\alpha_{t}}{\alpha_{t}-1}=1-\beta_{t}$. Since $\beta_{t}\in(0,1)$, $\alpha_{t}$ is monotonically decreasing with respect to $t$. We denote the quantized version of the noise estimation model as $\tilde{\mu}_{\theta}$. When the input remains unchanged, it is formulated by:
@@
where $\varepsilon_{t}$ means per-step quantization error which is introduced due to model quantization and only relies on the module at iteration $t$ and is independent of the others. The amount of error accumulated by continuously running the first Tt+1T-t+1 denoising steps is called the cumulative error $\delta_{t}$, then the input including cumulative error can be expressed as $\tilde{x}_{t}=x_{t}+\delta_{t}$. Consequently, referring to Eq. 3, the iterative process of the quantized model can be expressed as:
@@
Based on Eq. 4 and the definition of cumulative error, Eq. 5 can be reformulated as:
@@
Applying the first-order Taylor expansion, $\mu_{\theta}(x_{t}+\delta_{t},t)$ is approximated as $\mu_{\theta}(x_{t},t)+\mathbf{J}_{x_{t}}\delta_{t}$. Substituting it into Eq. 6, we can obtain the error propagation equation that relates the per-step quantization error to the cumulative error:
@@
in which $A_{t}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}I+B_{t}*\mathbf{J}_{x_{t}}$, $B_{t}=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}$ and $J_{x_{t}}=\nabla_{x_{t}}\mu_{\theta}(x_{t},t)$ is the Jacobian matrix of the denoising model $\mu_{\theta}$. Given that $\delta_{T}=0$, when we recursively expand Eq. 7 from $T$ to $t$, the cumulative error $\delta_{t}$ can be derived as:
@@ 3.2 Simplify Computational Complexity
Consequently, the inverse of the propagation matrix $\mathbf{A}_{j}$ can be approximately represented as $\frac{\sqrt{\alpha_{j}}}{\sqrt{\alpha_{j-1}}}\mathbf{I}$. By expanding the product terms in Eq. 8 and reducing the intermediate terms, the final result can be expressed as:
@@
Plugging Eq. 9 into Eq. 8, we have:
@@
Therefore, we can effectively reformulate Eq. 10 as:
@@
How to determine the value of the parameter $m$? Based on our discussion in Sec. 3.1, the cumulative error at step t1t-1 is related to the cumulative error $\delta_{t}$ from steps $T,\ldots,t+1]$ and the per-step quantization error $\varepsilon_{t}$ at step $t$. By substituting the correction term $\Delta_{t}$ into Eq. 7, we have:
@@
where $\widehat{\delta_{t-1}}$ is the corrected cumulative error and it should exhibits a strictly smaller upper bound in norm compared to $\delta_{t-1}$, signifying a more refined and accurate error representation. Based on this, we can solve for the reasonable value of $m$. Under mild regularity conditions, there exists $\sigma>0$ independent of the timestep $k$, such that the per-step quantization error satisfies $\|\varepsilon_{k}\|\leq\sigma,\forall k$. Based on this condition, we can deduce from $\|\widehat{\delta_{0}}\|\leq\sigma\sum_{t=1}^{T}\left(\prod_{k=1}^{t-1}\rho_{k}\right)\|\mathbf{B}_{t}\|\leq\|\delta_{0}\|$ that $m=1$. We provide a complete proof to this theorem in Appendix B and an empirical study demonstration in Appendix C. This implies that the cumulative error at any timestep is only related to the per-step quantization errors of the two immediately preceding steps. Thus, the final cumulative error correction term can be reformulated as: